# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v3

@@ I Introduction
Traditional GT and its performance (i.e., the tradeoff between the number of pool tests and recovery algorithm complexity), focused on the probabilistic model in which the items are identically distributed and independent [19, 14]. Recent research explores cases where prior information about the correlation of objects is available [20, 2, 3, 21, 22]. The motivation for this approach arises from the fact that correlated prior statistics have the potential to achieve higher recovery rates and reduce the number of required tests. In disease detection, leveraging information about the proximity between individuals, represented by contact tracing information or graphs, can lead to significant savings in pool tests [20, 2, 3, 21]. However, previous GT works presented solutions designed for specific models and applications and may not be extended easily to other models and applications. In numerous signal processing applications, correlation between different frequencies, time signals, or among different sensors can also be utilized to achieve more precise estimations [23, 13, 24]. Hidden Markov Model (HMM) is a common model for many physical signals, such as speech signals [25], human motion [26], and spectrum occupancy in communication systems [27]. Infections can be also modeled as HMMs [28]. For example, [29] presents a GT solution for a specific HMM derived from contact tracing. To the best of our knowledge, no existing solution addresses the GT problem with general Markovian priors and applicable to a wide range of diverse applications.
@@
In this work, we introduce a practical non-adaptive Multi-Stage GT (MSGT) algorithm for correlated items with prior statistics. The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests. The proposed algorithm can be applied for any statistical prior represented in trellis [33], e.g., finite-state machines and Markov processes. Using LVA, MSGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound. Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity. We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency. We also provide an analytical sufficiency bound to the number of pooled tests required by any maximum a-posteriori (MAP) decoder. This bound applies to any arbitrary underlying correlation between infected items. Our numerical results demonstrate that in practical regimes for COVID-19 [34, 35] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
@@ II Problem Formulation
Given a set of individuals $\mathcal{N}$, the objective in GT is to discover a small subset $\mathcal{K}$ of unknown defective items using the minimum number of measurements $T$. Let $N=|\mathcal{N}|$, $K=|\mathcal{K}|$ denote the total number of items and the number of defective items, respectively, where $K=\mathcal{O}(N^{\theta})$ for some $\theta\in\left[0,1\right)$. The binary vector $\mathbf{U}\in\left\{0,1\right\}^{N}$ represents the population, such that $U_{i}=1$ indicates that the $i$-th item is defective. We assume that the set of the individuals is sparse, such that $\theta\leq 1/3$[17, 19, 36, 14], and that each item can be in one of 2 states: defective and not-defective. For each item, there is some prior probability that it is defective, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\in\left[0,1\right]^{N}$, $\pi_{i}=P\left(U_{i}=1\right)$. The correlation between the state of the $i$-th item and the states of the previous $\tau$ items is represented by $\mathbf{\Phi}_{i}\in\left[0,1\right]^{2^{\tau}\times 2}$. $\tau=1,2,\dots$ represents the number of memory steps of the underlying process considered in the recovery stage. When there is no memory in the process, $\tau=0$ and $\{\pi_{i}\}_{i=1}^{N}$ hold all the prior information. $\mathbf{\Phi}_{i}[l,k]=P\left(U_{i}=s_{k}\lvert\left(U_{i-\tau},...,U_{i-1}\right)=s_{l}\right)$, where $s_{l}$ is the binary representation of $l$ by a $\tau$-length binary vector and $s_{k}\in\{0,1\}$ (For example see Fig. 6(b)).
@@
where $\bigvee$ is the Boolean OR function.
@@
To conclude, for a sparse subset of infected items of size $K$, form $N$, the goal in non-adaptive GT with correlated prior data items, is to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors, such that by observing $Y^{T}$ we can identify the subset of infected items with high probability and with feasible computational complexity. Thus, given the knowledge of the correlated prior data items and the available computational resources, the test designer could design the testing matrix and a recovery algorithm to maximize the success probability.
@@ III Main Results
In this section, we introduce the efficient multi-stage recovery algorithm for any statistical prior represented in a trellis diagram [33], detailed in Algorithm 1. In the first stage, standard low-complexity algorithms [19] reduce the search space independently of prior correlations. This reduction is guaranteed by new analytical results we derive. In the second stage, the algorithm employs a novel adaptation of the List Viterbi Algorithm (LVA) [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information. Additionally, we derive a bound to ensure the low complexity of the entire algorithm. Section III-A describes the proposed algorithm. Section III-B provides analytical results, followed by a discussion in Section  III-C. For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix A.
@@ III-A2 Recovery Process
In the second step of Stage 1, we use Definite Defectives (DD) algorithm [19] (line 3, Fig. 2.(b)), which goes over the testing matrix and the test result, looking for positive pool tests that include only one possibly defective item. DD denotes those items as definitely defective items and outputs them as the set $\mathcal{P}^{(DD)}$.
@@
In the first step of Stage 2, we translate the data we obtained in DND and DD, $\mathcal{P}^{(S_{1,2})}$ and $\mathcal{P}^{(DD)}$, into the state space in terms of transition probabilities, $\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$, and initial probabilities, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$, so we can employ all the gathered information in the next steps (line 6, Fig. 3.(a)). In the state space, the population sequence, $\mathbf{U}$, is parallel to time steps considered traditionally in VA for communications problems [32], and there are two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
@@ III-B Analytical Results
$P_{e,u}^{(DND)}$ $\leq P_{e,a}^{(DND)}$
$+\left(1+\frac{\left(\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)\right)^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1},$
@@
$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]=\mathbb{E}\left[P_{s}^{(DD)}\right]K\geq\\ \left(1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}\right)K.$
@@
We follow the lines of the success probability analysis in [36, Theorem 1]. Consider a list decoder that, given the test outcome $Y$, outputs $\mathcal{L}(Y)$, a list of items estimated as defective, such that $\lvert\mathcal{L}(Y)\lvert\geq K$. Let $N(Y)$ denote the number of possible defective sets for which, given $Y$, the list decoding is successful: $N(Y)\triangleq\sum_{\mathcal{K}}\mathbb{1}_{\{\mathcal{L}(Y)\cap\mathcal{K}=K\}}$. Assuming no prior information is given, $N(Y)$ is bounded by the counting bound
@@
It is important to note that unlike [36, Theorem 1], we halt the success probability analysis here without deriving a converse, and instead focus on the connection between $T$ and $\gamma$. This analysis is valid for any list decoder, particularly for an ML decoder [17, 44], allowing us to set $P_{s}=\eta$. Since MSGT uses MAP, the non-uniform and correlated prior information allows us to achieve better recovery performance for the same $T$ and $L$ compared to ML. Moreover, since for any $\gamma\geq 1$, $\gamma K$ is the maximum list size of estimated defective items we allow in MSGT. By substituting $L=\gamma K$, we establish a lower limit for $\gamma$ that guarantees MSGT’s success probability remains at least equal to that of the ML estimator. ∎
@@
In Stage 1, the complexity of DND is $\mathcal{O}\left(KN\log_{2}{N}\right)$ as analyzed in [45, Remark 6]. Then, for each positive entry of the test result vector $Y^{T}$, the DD algorithm counts the number of possibly defective items that participate in the corresponding pool test. That requires $K\left|\mathcal{P}^{(S_{1,1})}\right|\log_{2}{N}$ computations, and for simplification, we bound it by the DND complexity, i.e., $\mathcal{O}\left(KN\log_{2}N\right)$.
(a) $N=500$ and $K=3$.
(b) $N=1024$ and $K=8$.
(c) $N=10000$ and $K=13$.
In Stage 2, parallel LVA requires $L$ times more computations than the VA [32]. VA calculates all the possible transition probabilities for each step in the sequence. In GT, this sequence is the order’s items sequence, where with the suggested algorithm, it is enough to consider only the $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|$ items as the sequence steps. The optional states are basically either “non-defective” or “defective”, so there are four possible transitions in each step of the trellis proposed herein for GT. Nevertheless, this algorithm can be implemented to leverage additional memory to decide the state of each item based on the preceding $\tau$ items. Consequently, LVA takes $2^{2\tau}L\left|\mathcal{P}^{(S_{1,2})}\right|$ computations. For the average case, we use the expectation bounds from Theorems 1 and 3, such that:
$\mathbb{E}\left[\left|\mathcal{P}^{\left(S_{1}\right)}\right|\right]=\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]-\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]$
$\leq\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K.$
@@
$\begin{split}K&+\left(N-K\right)\left(P_{e,a}^{(DND)}+\right.\\ &\left.\left(1+\frac{g^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1}\right),\end{split}$ (7)
@@
Note that from Theorem 4 and Remark 1, it follows that the proposed MSGT algorithm performs $\mathcal{O}\left(\frac{1}{L}\left(\frac{N}{\gamma K}\right)^{K}\right)$, times fewer computational operations compared to MAP.
@@ III-C Discussion
To the best of our knowledge, MSGT is the first GT algorithm to effectively leverage Markovian prior statistics. Unlike numerous previous approaches, MSGT utilizes initial probabilities and transition matrices without necessitating specific adjustments for new use cases. The algorithm offers the flexibility to be fine-tuned to optimize its performance in accordance with the available computational resources and the number of tests, $T$. The simple reduction of the search space in Stage 1 enables MSGT to handle challenging regimes with a small number of tests. Stage 2, particularly the LVA step, contributes to its high success probability. Additionally, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [32]. It is important to note that, as explained in [32], achieving optimal results is ensured only with a very large $L$, inevitably leading to complexity equivalent to MAP’s. However, as we empirically demonstrate in the following section, results equivalent to MAP’s can be achieved with reasonable complexity. Moreover, it is shown that MSGT addresses practical regimes, e.g., in COVID-19 detection [46] ($T=48$ for $(N,K)=(384,4)$), in communications [47] ($(N,K)=(10^{5},6)$), and in GT quantizer [8] ($(N,K)=(1024,16)$ [48]).
Another aspect of novelty in this work is the integration of Viterbi Algorithm into the GT problem. In the context of Markovian priors, one can think of the population’s sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps. In that case, the selection of a Viterbi decoder becomes natural, offering an optimal and efficient decoding solution. However, the most likely sequence of items does not necessarily include $K$ defective items. Particularly, in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations. As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K$. To address this, we employ LVA, which produces a list of the $L$ most likely sequences, such that choosing an appropriate value for $L$ guarantees a successful recovery.
@@
(a) Number of computational operations.
(b) Correlation matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
(c) Long memory priors - exact recovery.

### v4

@@ I Introduction
Traditional GT and its performance (i.e., the tradeoff between the number of pool tests and recovery algorithm complexity), focuses on the probabilistic model in which the items are identically distributed and independent [19, 14]. Recent research explores cases where prior information about the correlation of objects is available [20, 2, 3, 21, 22]. The motivation for this approach arises from the fact that correlated prior statistics have the potential to achieve higher recovery rates and reduce the number of required tests. In disease detection, leveraging information about the proximity between individuals, represented by contact tracing information or graphs, can lead to significant savings in pool tests [20, 2, 3, 21]. However, previous GT works presented solutions designed for specific models and applications and may not be extended easily to other models and applications. In numerous signal processing applications, correlation between different frequencies, time signals, or among different sensors can also be utilized to achieve more precise estimations [23, 13, 24]. Hidden Markov Model (HMM) is a common model for many physical signals, such as speech signals [25], human motion [26], and spectrum occupancy in communication systems [27]. Infections can be also modeled as HMMs [28]. For example, [29] presents a GT solution for a specific HMM derived from contact tracing. To the best of our knowledge, no existing solution addresses the GT problem with general Markovian priors and applicable to a wide range of diverse applications.
@@
In this work, we introduce a practical non-adaptive Multi-Stage Group Testing (MSGT) algorithm for correlated items with prior statistics. The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests. The proposed algorithm can be applied for any statistical prior represented in trellis [33], e.g., finite-state machines and Markov processes. Using LVA, MSGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound. Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity. We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency. We provide an analytical sufficiency bound for the number of pooled tests needed by any maximum a-posteriori (MAP) decoder. This bound holds for any correlation between infected items. It also applies to multiple access communication systems, where $K$ out of $N$ users transmit information simultaneously. This is considered in [34, 35, 15] without prior statistical information. In contrast, existing analyses with prior statistics assume all $N$ users transmit at the same time [36, 37, 38, 39, 40, 41, 42]. Our numerical results demonstrate that in practical regimes for COVID-19 [43, 44] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
@@ II Problem Formulation
Given a set of individuals $\mathcal{N}$, the objective in GT is to discover a small subset $\mathcal{K}$ of unknown defective items using the minimum number of measurements $T$. Let $N=|\mathcal{N}|$, $K=|\mathcal{K}|$ denote the total number of items and the number of defective items, respectively, where $K=\mathcal{O}(N^{\theta})$ for some $\theta\in\left[0,1\right)$. The binary vector $\mathbf{U}\in\left\{0,1\right\}^{N}$ represents the population, such that $U_{i}=1$ indicates that the $i$-th item is defective. We assume that the set of the individuals is sparse, such that $\theta\leq 1/3$[17, 19, 45, 14], and that each item can be in one of 2 states: defective and not-defective. For each item, there is some prior probability that it is defective, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\in\left[0,1\right]^{N}$, $\pi_{i}=P\left(U_{i}=1\right)$. The correlation between the state of the $i$-th item and the states of the previous $\tau$ items is represented by $\mathbf{\Phi}_{i}\in\left[0,1\right]^{2^{\tau}\times 2}$. $\tau=1,2,\dots$ represents the number of memory steps of the underlying process considered in the recovery stage. When there is no memory in the process, $\tau=0$ and $\{\pi_{i}\}_{i=1}^{N}$ hold all the prior information. $\mathbf{\Phi}_{i}[l,k]=P\left(U_{i}=s_{k}\lvert\left(U_{i-\tau},...,U_{i-1}\right)=s_{l}\right)$, where $s_{l}$ is the binary representation of $l$ by a $\tau$-length binary vector and $s_{k}\in\{0,1\}$ (for example, see Fig. 7(b)).
@@
where $\bigvee$ is the Boolean OR function. All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise. These definitions are illustrated in Fig. 1.
@@
In conclusion, the goal of non-adaptive GT with correlated prior data is to identify the $K$ defective items out of the $N$ items, where there is correlation between different infected items. To achieve this, one needs to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors. The algorithm can identify the subset of infected items by observing $Y^{T}$ with high probability and with feasible computational complexity. Given the knowledge of the correlated prior data items and the available computational resources, one can design the testing matrix and a recovery algorithm to maximize the success probability.
@@ III Main Results
In this section, we introduce the efficient multi-stage recovery algorithm for any statistical prior represented in a trellis diagram [33], detailed in Algorithm 1. In the first stage, standard low-complexity algorithms [19] reduce the search space independently of prior correlations. This reduction is guaranteed by new analytical results we derive. In the second stage, the algorithm employs a novel adaptation of the LVA [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information. Additionally, we derive a bound to ensure the low complexity of the entire algorithm. Section III-A describes the proposed algorithm. Section III-B provides analytical results, followed by a discussion in Section  III-C. For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix A.
@@ III-A2 Recovery Process
In the second step of Stage 1, we use the Definitely Defectives (DD) algorithm [19] (line 3, Fig. 2.(b)), which goes over the testing matrix and the test result, looking for positive pool tests that include only one possibly defective item. DD denotes those items as definitely defective items and outputs them as the set $\mathcal{P}^{(DD)}$.
@@
In the first step of Stage 2, we translate the data we obtained in DND and DD, $\mathcal{P}^{(S_{1,2})}$ and $\mathcal{P}^{(DD)}$, into the state space in terms of transition probabilities, $\{\mathbf{\Phi}_{i}\}_{i=1}^{N}$, and initial probabilities, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}$, so we can employ all the gathered information in the next steps (line 6, Fig. 3.(a)). In the state space, the population sequence, $\mathbf{U}$, is parallel to time steps considered traditionally in Viterbi Algorithm (VA) for communications problems [32], and there are two possible states per item, the first indicates “non-defective” and the second indicates “defective”.
@@ III-B Analytical Results
$P_{e,u}^{(DND)}$ $\leq P_{e,a}^{(DND)}+\left(1+\frac{\left(\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)\right)^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1},$
@@
$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]=\mathbb{E}\left[P_{s}^{(DD)}\right]K\geq\left(1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}\right)K.$
@@
We follow the lines of the success probability analysis in [45, Theorem 1]. Consider a list decoder that, given the test outcome $Y$, outputs $\mathcal{L}(Y)$, a list of items estimated as defective, such that $\lvert\mathcal{L}(Y)\lvert\geq K$. Let $N(Y)$ denote the number of possible defective sets for which, given $Y$, the list decoding is successful: $N(Y)\triangleq\sum_{\mathcal{K}}\mathbb{1}_{\{\mathcal{L}(Y)\cap\mathcal{K}=K\}}$. Assuming no prior information is given, $N(Y)$ is bounded by the counting bound
@@
It is important to note that unlike [45, Theorem 1], we halt the success probability analysis here without deriving a converse, and instead focus on the connection between $T$ and $\gamma$. This analysis is valid for any list decoder, particularly for an ML decoder [17, 53], allowing us to set $P_{s}=\eta$. Since MSGT uses MAP, the non-uniform and correlated prior information allows us to achieve better recovery performance for the same $T$ and $L$ compared to ML [^2] In Section IV-C, we provide sufficiency bound for any MAP decoder with an arbitrary underlying correlation between infected items that outperform the ML sufficiency bound and illustrate it for practical scenarios in Section V.. Moreover, since for any $\gamma\geq 1$, $\gamma K$ is the maximum list size of estimated defective items we allow in MSGT. By substituting $L=\gamma K$, we establish a lower bound for $\gamma$ that guarantees MSGT’s success probability remains at least equal to that of the ML estimator. ∎
@@
In Stage 1, the complexity of DND is $\mathcal{O}\left(KN\log_{2}{N}\right)$ as analyzed in [54, Remark 6]. Then, for each positive entry of the test result vector $Y^{T}$, the DD algorithm counts the number of possibly defective items that participate in the corresponding pool test. That requires $K\left|\mathcal{P}^{(S_{1,1})}\right|\log_{2}{N}$ computations, and for simplification, we bound it by the DND complexity, i.e., $\mathcal{O}\left(KN\log_{2}N\right)$.
In Stage 2, parallel LVA requires $L$ times more computations than the VA [32]. The VA calculates all the possible transition probabilities for each step in the sequence. In GT, this sequence is the order’s items sequence, where with the suggested algorithm, it is enough to consider only the $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|$ items as the sequence steps. The optional states are basically either “non-defective” or “defective”, so there are four possible transitions in each step of the trellis proposed herein for GT. Nevertheless, this algorithm can be implemented to leverage additional memory to decide the state of each item based on the preceding $\tau$ items. Consequently, LVA takes $2^{2\tau}L\left|\mathcal{P}^{(S_{1,2})}\right|$ computations. For the average case, we use the expectation bounds from Theorems 1 and 3, such that:
$\mathbb{E}\left[\left|\mathcal{P}^{\left(S_{1}\right)}\right|\right]=\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]-\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]\leq\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K.$
@@
$\begin{split}K&+\left(N-K\right)\left(P_{e,a}^{(DND)}+\left(1+\frac{g^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1}\right),\end{split}$ (7)
@@
Note that from Theorem 4 and Remark 1, it follows that the proposed MSGT algorithm performs $\mathcal{O}\left(\frac{1}{L}\left(\frac{N}{\gamma K}\right)^{K}\right)$ times fewer computational operations compared to MAP.
@@ III-C Discussion
To the best of our knowledge, MSGT is the first GT algorithm to effectively leverage Markovian prior statistics. Unlike numerous previous approaches, MSGT utilizes initial probabilities and transition matrices without necessitating specific adjustments for new use cases. The algorithm offers the flexibility to be fine-tuned to optimize its performance in accordance with the available computational resources and the number of tests, $T$. The simple reduction of the search space in Stage 1 enables MSGT to handle challenging regimes with a small number of tests. Stage 2, and the LVA step in particular, contributes to its high success probability. Furthermore, using the parallel implementation of LVA, rather than the iterative one, keeps the complexity low [32]. It is important to note that, as explained in [32], achieving optimal results is ensured only with a very large $L$, inevitably leading to complexity equivalent to MAP’s. However, as we empirically demonstrate in the following section, results equivalent to MAP’s can be achieved with reasonable complexity. Moreover, it is shown that MSGT addresses practical regimes, e.g., in COVID-19 detection [55] ($T=48$ for $(N,K)=(384,4)$), in communications [56] ($(N,K)=(10^{5},6)$), and in GT quantizer [8] ($(N,K)=(1024,16)$ [57]).
Another aspect of novelty in this work is the integration of the VA into the GT problem. In the context of Markovian priors, one can think of the population’s sequence of items as a sequence of observations stemming from a hidden Markov process within a given Markov model over $N$ steps. In that case, the selection of a Viterbi decoder becomes natural, offering an optimal and efficient decoding solution. However, the most likely sequence of items does not necessarily include $K$ defective items. Particularly in sparse signal scenarios, which is the focus of GT, the most likely sequence typically involves the minimum number of defectives that explain the observations. As a result, VA may not necessarily detect more defective items than already known and may detect even more than $K$. To address this, we employ LVA, which produces a list of the $L$ most likely sequences, such that choosing an appropriate value for $L$ guarantees a successful recovery.